OpenStudy (anonymous):

Simplify: (4a – 2b)(a + b)

OpenStudy (anonymous):

i believe you use FOIL...

OpenStudy (anonymous):

(4a-2b)(a+b) =(4a*a + 4a*b + (-2b)*a + (-2b)*b) =(4a² + 4ab - 2ba -2b²) =4a² + 2ab - b²

jimthompson5910 (jim_thompson5910):

you can also use this rectangle